OpenStudy (anonymous):

Given the trinomial 3x^2 - 6x + 5, what is the value of the discriminant?

OpenStudy (anonymous):

discriminant =\[b ^{2}-4ac\]

OpenStudy (anonymous):

when you plug in the values you get b=-6 b^2=36 -4ac=-60

OpenStudy (charlotte123):

@croslali good job :)

OpenStudy (anonymous):

thank you! i'm so confused about all this trinomial stuff.

OpenStudy (anonymous):

and then you get discriminant =36-60=-14

OpenStudy (anonymous):

@14whitea i see that you are new here may you enjoy your experience at openstudy have a nice day dont forget to fan me from my homepage #rawr :3

OpenStudy (anonymous):

welcome ^_^ since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me. Have fun learning with Open Study! :)